40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 13 21:09:03 2013, manasa, Update, Green Locking, Arms ready for ALS 
    Reply  Thu Aug 15 12:30:17 2013, manasa, Update, Green Locking, Lost beat notes 
       Reply  Thu Aug 15 19:05:07 2013, manasa, Update, Green Locking, ALS out of loop noise ALS_OutLoop.pdf
Message ID: 9015     Entry time: Thu Aug 15 19:05:07 2013     In reply to: 9014
Author: manasa 
Type: Update 
Category: Green Locking 
Subject: ALS out of loop noise 

Beat notes were recovered for both the arms.

I locked the arms to IR using PDH and measured the ALS out of loop noise at the phase tracker output.

The Y arm has the same 300Hz/rtHz rms. The X arm rms noise measures nearly the same as the Y arm in the 5-500Hz region (X arm has improved nearly 10 times after the last whitening filter stage change  old elog ).

The noise in the ALSX error signals could be related to the bad alignment and conditions at the X end.

Attachment 1: ALS_OutLoop.pdf  70 kB  Uploaded Thu Aug 15 20:05:25 2013  | Hide | Hide all
ALS_OutLoop.pdf
ELOG V3.1.3-